OpenStudy (anonymous):

hear is the question Determine the sum of the following geometric series. 1/64+1/16+1/4+...+1024

OpenStudy (anonymous):

2^-6+...+2^10 The formula for geometric series: http://en.wikipedia.org/wiki/Geometric_progression#Geometric_series To shift the sequence, change 'a'

OpenStudy (mathmath333):

\(\large\tt \color{black}{\dfrac{a(r^n-1)}{r-1}}\) \(\large\tt \color{black}{\text{where a = first term and r is common ratio}}\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you know how to find the common ratio

OpenStudy (mathmath333):

its \[\frac{ 1 }{ 16 }\div \frac{ 1 }{ 64}\] its \(\large\tt \color{black}{\text{2nd term divided by first term or }}\) \(\large\tt \color{black}{\text{3rd term divided by 2nd term or }}\) \(\large\tt \color{black}{\text{4th term divided by 3rd term }}\)

OpenStudy (anonymous):

oh okay

OpenStudy (mathmath333):

\(\large\tt \color{black}{\text{and n = number of terms in series}}\)